SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (May 23, 2006)

Attendees: Paul Dauncey, Al Eisner, Fang Fang, Arik Kreisel, Selina Li, Dan Walker

Operations

  • After the ROD last week, PEP recovered quickly with luminosity increasing. Now We reached all-time record for luminosity (10.028 x 10^33).
  • Al suggested that it's time to take a physics_l1passthru run for studies such as trigger rate and behavior of L3, etc. Selina made a request at the 3:45 operation meeting, but we will wait till Rainer confirms that the configuration key is up-to-dated.

EMT

  • Paul would like to know what the impact of Chris O'Grady's deadtime bug fix is. In the past week weeks or so, the running condition wasn't stable, so it might be hard to tell. But lately with good running condition, we should check again to see if there was a dramatic improvement or not.

DCT/DCZ

  • Dan and Selina tested a new version of the firmware last Tuesday (May 16th), but it caused a lot of ZPD damages. So last Wednesday (during the ROD), we put the old TSF firmware back to the TSF-X and TSF-Y crates. There is still one occurance of out-of-order and out-of-synch errors (this morning owl shift) for about 2 seconds since then. These errors are similar to the one we got from the versions we had been using, namely the first 8 slots of the boards showed out-of-order from the rest of the slots in TSF-X.

L3

  • Arik would like to get some help with C++ coding in general and in particular when reading the L3 tracking codes. If you are a C++ expert and willing to help, you can contact him so he can go to you for help when needed.

Data Quality and Monitoring

  • Selina made a simple modification (Changed "Fast_Recent" to "Medium_Full" for the scaler group) so that we get about 45 minutes of the run rather than only 15 minutes for the time history plots in the ps archive. The new tags for L1TFMon are V00-05-16 for release 18 and V00-06-05 for release 19. This online release hasn't been installed at IR2 yet.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.